PRECISE THEORETICAL STUDY OF SPECTROSCOPIC CONSTANTS IN DIATOMICS

\underline{L.V. SKRIPNIKOV}, A.N. PETROV, A.V. TITOV, N.S. MOSYAGIN^{a},

B.P. Konstantinov Petersburg Nuclear Physics Institute, Gatchina, Leningrad district 188300; Department of Physics, Saint Petersburg State University, Petrodvoretz 198904, RUSSIA.

During the last decade a significant progress in methods of molecular spectroscopy was achieved. One of the most important applications of these methods is a measurement of an electron electric dipole moment in diatomic molecules and cations containing heavy elements. In order to speed up such experiments, which usually prepared for many years, one can apply accurate ab initio quantum-chemical methods to predict different spectroscopic constants of a molecule or cation under consideration: scheme of electronic terms, vibrational and rotation constants, hyperfine structure constants, g-factors, lifetimes etc. In such calculations both correlation and relativistic effects should be taken into account. This is not a trivial problem for systems containing transition elements and especially lanthanides and actinides. We report results of our recent investigations of a number of diatomics including theoretical investigation of HfF^{+} ^{b}. Details of used methods are discussed.

aSPbU Fundamental Science Research grant from Federal budget N 0.038.652.2013 and RFBR grant 13-02-01406 are gratefully acknowledged

bK. C. Cossel, D. N. Gresh, L. C. Sinclair, T. Coffey, L. V. Skripnikov, A. N. Petrov, N. S. Mosyagin, A. V. Titov, R. W. Field, E. R. Meyer, E. A. Cornell and J. Ye, Chem. Phys. Lett., 546, 1 (2012).